OpenStudy (anonymous):

let f(x)-4x+7 and g(x)=10x-6, find f(g(x))

OpenStudy (ranga):

f(x) = 4x + 7 ----- (1) g(x) =10x - 6 ----- (2) f(g(x)) = f(10x-6) On the RHS of (1), replace all x with (10x-6) and simplify to get f(g(x)).